# Electromagnetic Follow-up to Gravitational Wave Events with the UltraViolet EXplorer (UVEX)

### v2

@@ 6.3 Assumed Intrinsic KN Magnitude
Perhaps unsurprisingly, this strategy carries the most significant effect on the number of UVEX EM-GW ToO triggers, almost doubling the ToO trigger count compared to the fiducial strategy. The M-13.1 variation alone results in an estimated $6.8^{+8.3}_{-3.8}$ and $11.2^{+13.9}_{-6.3}$ additional ToO triggers for O5 and O6, for a total of $15.7^{+19.4}_{-8.8}$ and $22.5^{+27.9}_{-12.6}$ triggers, respectively. We also consider a joint strategy, dubbed M-13.1 + CT90, which couples M-13.1 with a relaxed coverage threshold of 90%. This approach yields an estimated $15.0^{+18.6}_{-8.4}$ and $21.8^{+27.0}_{-12.2}$ additional triggers for O5 and O6, for a total of $23.9^{+29.7}_{-13.4}$ and $33.1^{+41.0}_{-18.5}$ triggers, respectively. By and large, the additional triggers result from the increased distances at which this strategy will trigger ToOs, as can be seen in Fig. 4 as compared to Figs. 2 and 3. While adopting the M-13.1 + CT90 strategy does result in a substantially reduced success rate (63.5% and 63.6% for the O5 and O6 scenarios), it nonetheless yields the highest estimate for successful counterpart detections: $15.2^{+18.8}_{-8.6}$ ($21.1^{+26.1}_{-11.9}$) for O5 (O6). The trigger counts, success rates, and estimated counterpart detections for each strategy can be found in Table 3.
Figure 4: UVEX ToO selection for the O6 simulation with a 90% threshold for minimum coverage by UVEX of the GW localization region and an assumed intrinsic KN magnitude of -13.1 AB mags. Points correspond to individual simulated events from the Kiendrebeogo et al. (2023) O6 observing scenario, so the total number of points is not one-to-one with the expected number of observed BNS mergers (although it is proportional); scaling from the simulated event rate to an astrophysical BNS rate of $210^{+240}_{-120}\ \mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ (Abbott et al., 2023) yields $14.7^{+18.3}_{-8.3}$ UVEX ToO triggers over 1.5 years of O6. Events are plotted by their mean GW distance estimate and 90% C.L. GW localization areas. Points in grey are not selected due to the overall threshold of $\leq 100$ deg2 GW 90% C.L. localization area. Events with $\leq 100$ deg2 localizations that are not selected due falling outside of the UVEX FoR or exceeding the maximum epoch time constraint of 3 hrs are marked as black and orange $\times$×s, respectively. Selected events are color-coded by per-tile exposure time.

### v4

@@ 6.3 Assumed Intrinsic KN Magnitude
Perhaps unsurprisingly, this strategy carries the most significant effect on the number of UVEX EM-GW ToO triggers, almost doubling the ToO trigger count compared to the fiducial strategy. The M-13.1 variation alone results in an estimated $6.8^{+8.3}_{-3.8}$ and $11.2^{+13.9}_{-6.3}$ additional ToO triggers for O5 and O6, for a total of $15.7^{+19.4}_{-8.8}$ and $22.5^{+27.9}_{-12.6}$ triggers, respectively. We also consider a joint strategy, dubbed M-13.1 + CT90, which couples M-13.1 with a relaxed coverage threshold of 90%. This approach yields an estimated $15.0^{+18.6}_{-8.4}$ and $21.8^{+27.0}_{-12.2}$ additional triggers for O5 and O6, for a total of $23.9^{+29.7}_{-13.4}$ and $33.1^{+41.0}_{-18.5}$ triggers, respectively. By and large, the additional triggers result from the increased distances at which this strategy will trigger ToOs, as can be seen in Fig. 1. While adopting the M-13.1 + CT90 strategy does result in a substantially reduced success rate (63.5% and 63.6% for the O5 and O6 scenarios), it nonetheless yields the highest estimate for successful counterpart detections: $15.2^{+18.8}_{-8.6}$ ($21.1^{+26.1}_{-11.9}$) for O5 (O6). The trigger counts, success rates, and estimated counterpart detections for each strategy can be found in Table 4.
Figure 1: UVEX ToO selection summary plots for selected simulations discussed in §56. Top: event selection for the fiducial strategy (§5) as applied to the O5 and O6 observing scenarios. Bottom: event selection for selected expanded ToO strategies. The bottom left panel shows the CT90 strategy (§6.2), which reduces the minimum required UVEX coverage of the GW localization region to 90%. The bottom right panel shows the M-13.1+CT90 strategy (§6.3), which additionally assumes a less-conservative value of the intrinsic KN luminosity for exposure time calculations. Points correspond to individual simulated events from the Kiendrebeogo et al. (2023) observing scenarios, so the total number of points is not one-to-one with the expected number of observed BNS mergers (although it is proportional); scaling from the simulated event rate to an astrophysical BNS rate of $210^{+240}_{-120}\ \mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ (Abbott et al., 2023) yields the UVEX BNS ToO trigger rates given in Table 4. Events are plotted by their mean GW distance estimate and 90% C.L. GW localization areas. Points in grey are not selected due to the overall threshold of $\leq 100$ deg2 GW 90% C.L. localization area. Events with $\leq 100$ deg2 localizations that are not selected due falling outside of the UVEX FoR or exceeding the maximum epoch time constraint of 3 hrs are marked as black and orange $\times$×s, respectively. Selected events are color-coded by per-tile exposure time.